OpenStudy (anonymous):

Suppose that the functions q and r are defined as follows. q(x)=x^2+5 r(x)= Square Root of X+7 Find the following: (r o q)(2) = ? (q o r)(2) = ?

OpenStudy (anonymous):

so (r o q)(2) is a fancy way of writing r(q(2)) which means what do you get when you plug 2 into q and then that answer into r

OpenStudy (anonymous):

so for the first one we plug 2 into q 2^2+5 4+5 9

OpenStudy (anonymous):

now we plug that into r 9+7 15\[\sqrt{15} \]

OpenStudy (anonymous):

So then how do you differentiate (r o q) and (q o r)? I guess more than anything I'm just confused on how to go about come about the answer. So do you think plug in \[\sqrt{15}\](9(2)) to solve the first part, then reverse for the second? Or am I missing something. Appreciate the help.

OpenStudy (anonymous):

wait so what is the original question you are asked then. Is it just function composition or is it asking you to differentiate the composition?

OpenStudy (anonymous):

because above it looks to be that all your asked to do is find the composition of the functions

OpenStudy (anonymous):

oh sorry you mean DIFFERENTIATE as in tell the difference sorry I thought you meant differentiate as in take the derivative

OpenStudy (anonymous):

ok the way I like to think of it is like an onion

OpenStudy (anonymous):

you work from the inside out or in the case above right to left

OpenStudy (anonymous):

The question I need to answer is at the bottom, find the following, and plug in the answers where the question marks are. Is what you showed me all I need to do? It makes sense to me how you got there I just feel like I'm missing something to show two different answers

OpenStudy (anonymous):

on the far right is 2

OpenStudy (anonymous):

then to the left of that is q so you take 2 and plug it into q

OpenStudy (anonymous):

then to its left is r so then you plug your answer from q into r

OpenStudy (anonymous):

with the second part you see that r is to the left of 2 so you take 2 plug it into r and then into q since q is to the left of r

OpenStudy (anonymous):

alright give me a sec to see if I understand that

OpenStudy (anonymous):

r=\[\sqrt{2+7}\] which simplifies into r=3 and then q=3^2+5= 9+5=14 q=14 So then for the (r o q)(2) q=2^2+5=9 and \[\sqrt{9+7}\] = \[\sqrt{16}\] so r=4 Then what's the final step to find the solution? Or is it done? I feel like I'm making it harder than this needs to be haha

OpenStudy (anonymous):

Would both just simply = 4?

OpenStudy (anonymous):

Whoops, added an extra step on that

OpenStudy (anonymous):

Wait no I didn't, forgot I did both haha. Sorry, my brain is melted it's been a long day

OpenStudy (anonymous):

That's all there is you just plug it in like that and I don't know what you mean by them both simplifying to four. the second part with sqrt(16) obviously simplifies to four but q=14 would be the end of it

OpenStudy (anonymous):

and sorry for my above mistake of 9+7=15. I face palmed at that one. You are right it is 16 and then the sqaure root of 16

OpenStudy (anonymous):

Yep I realized that mistake, so (r o q)(2)=4, and (q o r)(2)=14? Thanks a lot for your help, that onion analogy really helped. I was making things more complicated than they needed to be